(5) The half life for the decomposition of N2O5 in troposphere is 2.05 x 104 s. How long will
it take for a sample of N2O5 to decay to 60% of its initial value?
1
Expert's answer
2020-12-03T14:30:05-0500

Solution.

"t(1/2) = \frac{ln(2)}{k}"

"k = \frac{ln(2)}{t(1/2)}"

"k = \frac{1}{t} ln(C0/C)"

"t = \frac{ln(\frac{C0}{C})}{k}"

C0 = 1

C = 0.6*C0

"k = 3.38 \times 10^{-5} sec"

t = 15113 sec = 4.2 hours

Answer:

t = 4.2 hours